# Sampling from the Mean-Field Stationary Distribution

### v2

@@ 4.2 General McKean–Vlasov Setting
In the general setting, under Assumptions 56, and 7, the work of [SWN23] provided the first discretization bounds. They impose further assumptions and their resulting complexity bound is rather complicated, but it reads roughly $MN=\widetilde{\mathcal{O}}(\nicefrac{{\operatorname{poly}(\kappa)\,d^{2}}}{{\varepsilon^{4}}})$ for the discretization of ($\mathsf{gMV}_{N}$). Subsequently, [FW23] obtained an improved complexity of $MN=\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{4}d^{3/2}}}{{\varepsilon^{3}}})$ via ULMC in the averaged $\mathsf{TV}$ distance. In comparison, we can improve this complexity guarantee to $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{5/2}d^{3/2}}}{{\varepsilon^{3}}})$, and the guarantee even holds in $\sqrt{\operatorname{\mathsf{KL}}}$ if we combine ULMC with the proximal sampler. It appears that we gain one factor of $\sqrt{\kappa}$ through sharper discretization analysis (via [Zha+23], or via the error analysis of the proximal sampler in [AC23]), and one factor of $\kappa$ via a sharper propagation of chaos result (Theorem 5).
We also note that the result of [FW23] is based on a kinetic version of the propagation of chaos argument from [Che+23], whereas our approach uses the original “non-kinetic” argument from [CRW22] in the form of Theorem 5.
@@
Let us consider the problem of learning a two-layer neural network in the mean-field regime. Let $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ be a function parameterized by $\theta\in\mathbb{R}^{p}$, and for any probability measure $\mu$ over $\mathbb{R}^{p}$, let $f_{\mu}\coloneqq\int f_{\theta}\,\mu(\mathrm{d}\theta)$. For example, in a standard two-layer neural network, we take $\theta=(a,w)\in\mathbb{R}\times\mathbb{R}^{d}$ and $f_{a,w}(x)=a\,\mathsf{ReLU}(\langle w,x\rangle)$. When $\mu=\frac{1}{m}\sum_{j=1}^{m}\delta_{(a_{j},w_{j})}$ is an empirical measure, then $f_{\mu}$ is the function computed by a two-layer neural network with $m$ hidden neurons. In this formulation, however, we can take $\mu$ to be any probability measure, corresponding to the mean-field limit $m\to\infty$ [CB18, MMN18, Chi22, RV22, SS20].
@@
Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ in $\mathbb{R}^{d}\times\mathbb{R}$ and a loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, we can formulate neural network training as the problem of minimizing the loss $\mu\mapsto\sum_{i=1}^{n}\ell(f_{\mu}(x_{i}),y_{i})$. To place this within the general McKean–Vlasov framework, we add two regularization terms: (1) $\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$ corresponds to weight decay; and (2) $\frac{\sigma^{2}}{2}\int\log\mu\,\mathrm{d}\mu$ is entropic regularization. We are now in the setting of §2.1.2, with $\mathcal{F}_{0}(\mu)=\sum_{i=1}^{n}\ell(f_{\mu}(x_{i}),y_{i})$. To minimize this energy, it is natural to consider the Euler–Maruyama discretization of ($\mathsf{gMV}_{N}$), which corresponds to learning the neural network via noisy GD, and was considered in [SWN23]. Recent works [Che+23, FW23] also considered the underdamped version of ($\mathsf{gMV}$) and its discretization. Under the assumptions common to those works as well as our own, our results yield improved algorithmic guarantees for this task (see Example 13).
@@
Unfortunately, the assumptions used for the analysis of the general McKean–Vlasov are restrictive and limit the applicability to neural network training. For example, it suffices for $\ell$ to be convex in its first argument (to satisfy Assumption 5), to have two bounded derivatives (w.r.t. its first argument), and for $\theta\mapsto f_{\theta}(x_{i})$ to have two bounded derivatives for each $x_{i}$. The last condition is satisfied, e.g., for $f_{\theta}(x)=\tanh(\langle\theta,x\rangle)$. For a genuinely two-layer example, we can take $f_{\theta}(x)=\tanh(a)\tanh(\langle w,x\rangle)$ for $\theta=(a,w)\in\mathbb{R}\times\mathbb{R}^{d}$. Under these conditions, Assumptions 6 and 8 hold, which in turn furnish log-Sobolev inequalities via Lemmas 6 and 9. In general, these LSI constants depend exponentially on quantities such as the dimension $p$ of the parameter space, which is unavoidable as global non-convex optimization is intractable in the worst case. However, we note that these limitations are inherited from the prior literature and not specific to our approach.
@@ 5 Conclusion
We conclude by listing some future directions of study. We believe there is further room for improvement in the propagation of chaos results. For example, can the sharp rate in Theorem 3 be extended to stronger metrics such as Rényi divergences, as well as to situations when the weak interaction condition (Assumption 3) fails, e.g., in the strongly displacement convex case or in the setting of §3.1.2? For the sampling guarantees, the prior works [BS23, SWN23] considered different settings, such as potentials satisfying convexity at infinity or the use of stochastic gradients; these extensions are compatible with our approach and could possibly lead to improvements in these cases, as well as others. Finally, it would be interesting to develop further applications of sampling from the mean-field distribution, e.g., applications to neural network training under less stringent hypotheses.
@@ Acknowledgements
YK was supported in part by NSF awards CCF-2007443 and CCF-2134105. MSZ was supported by NSERC through the CGS-D program. SC was supported by the Eric and Wendy Schmidt Fund at the Institute for Advanced Study. MAE was supported by NSERC Grant [2019-06167] and CIFAR AI Chairs program at the Vector Institute. MBL was supported by NSF grant DMS-2133806.
range
pages8 rangepages-1 rangepages54 rangepages29 rangepages18 rangepages88 rangepages36 rangepages-1 rangepages51 rangepages28 rangepages121 rangepages157 rangepages31 rangepages2 rangepages41 rangepages26 rangepages-1 rangepages32 rangepages-1 rangepages49 rangepages18 rangepages44 rangepages20 rangepages11 rangepages36 rangepages24 rangepages69 rangepages17 rangepages37 rangepages38 rangepages32 rangepages58 rangepages35 rangepages24 rangepages21 rangepages5 rangepages-1 rangepages54 rangepages17 rangepages17 rangepages74 rangepages40 rangepages49 rangepages47 rangepages28 rangepages87 rangepages49 rangepages235 rangepages-1 rangepages935 rangepages63 rangepages16 rangepages36
@@ Appendix A Control of the Finite-Particle Error
Lemma 14 (Talagrand’s Transport Inequality, [OV00]).
@@ A.1 LSI Case
$-\frac{\sigma^{2}}{2}\,\nabla_{x^{i}}\log\mu^{1:k}(x^{1:k})=\nabla V(x^{i})+\frac{1}{N-1}\sum_{\begin{subarray}{c}j=1\\ i\neq j\end{subarray}}^{k}\nabla W(x^{i}-x^{j})+\frac{N-k}{N-1}\operatorname{\mathbb{E}}_{\mu^{k+1|1:k}(\cdot\mid x^{1:k})}\nabla W(x^{i}-\cdot)\,.$
@@ A.1.1 Bounding the Error Terms
$\qquad+(k-1)\,(k-2)\operatorname{\mathbb{E}}\bigl{\langle}\nabla W(X^{1}-X^{2})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot),W(X^{1}-X^{3})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot)\bigr{\rangle}$
@@
$\qquad+(k-1)^{2}\operatorname{\mathbb{E}}\bigl{\langle}\nabla W(X^{1}-X^{2})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot),\,\nabla W(X^{1}-X^{3})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot)\bigr{\rangle}\,,$
@@
where the second inequality follows from (TI), and the last one follows from the data-processing inequality for the $\operatorname{\mathsf{KL}}$ divergence.
For the second term, the Cauchy–Schwarz inequality leads to
@@ A.2 Strongly Convex Case
The following propagation of chaos argument for the strongly log-concave case is based on [Szn91]. Let $(X^{1:N}_{t})_{t\geq 0}$ denote the stochastic process following the finite-particle stochastic differential equation ($\mathsf{pMV}_{N}$). Let the corresponding semigroup be denoted ${(\mathcal{T}_{t})}_{t\geq 0}$, defined as follows. For any test function $f:\mathbb{R}^{d\times N}\to\mathbb{R}$,
@@
Proof.  We resort to a coupling argument, noting that $\pi$ is stationary under ($\mathsf{pMV}$). Starting with $\pi^{\otimes N}$, we evolve ${(X^{1:N}_{t})}_{t\geq 0}$ and ${(Y^{1:N}_{t})}_{t\geq 0}$ according to ($\mathsf{pMV}_{N}$) and ($\mathsf{pMV}$) respectively, i.e., $X^{1:N}_{t}\sim\pi^{\otimes N}\mathcal{T}_{t}$ and $Y^{1:N}_{t}\sim\pi^{\otimes N}$. This argument is adapted from the original propagation of chaos proof by [Szn91].
@@ A.3 General Functional Case
For any measure $\mu$, define its entropy as $\operatorname{\mathsf{ent}}(\mu)=\int\log\mu\,\mathrm{d}\mu$. We now provide a self-contained propagation of chaos argument in the general McKean–Vlasov setting, following [CRW22]. We begin with the following entropy toast inequality, i.e., half of the entropy sandwich inequality from [CRW22].
@@
$\mathcal{E}^{N}(\nu^{1:N})=N\int\mathcal{F}(\rho_{x^{1:N}})\,\nu^{1:N}(\mathrm{d}x^{1:N})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\mu)\,.$
@@
[…]
@@
$\qquad=-\frac{\sigma^{2}}{2}\int\sum_{i=1}^{N}\log\pi(x^{i})\,\nu^{1:N}(\mathrm{d}x^{1:N})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\nu^{1:N})\,,$
where (i) follows from that $\pi$ is the minimizer of $\mathcal{E}$. The last line is just $\frac{\sigma^{2}}{2}\operatorname{\mathsf{KL}}(\nu^{1:N}\mathbin{\|}\pi^{\otimes N})$, so we are done. ∎
@@
In order to relate the first term $\mathsf{A}$ to the $\operatorname{\mathsf{KL}}$ divergence, for each $x^{-i}\in\mathbb{R}^{d\times(N-1)}$ we introduce the probability measure $\tau_{x^{-i}}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ via
@@
$=\log\int\exp\Bigl{(}-\log\pi(z)-\frac{2}{\sigma^{2}}\,\delta\mathcal{F}(\rho_{x^{-i}},z)\Bigr{)}\,\pi(\mathrm{d}z)$
@@
$\zeta_{x^{1:N}}^{i}(y)\coloneqq\delta\mathcal{F}(\rho_{x^{-i}},y)-\delta\mathcal{F}(\rho_{x^{1:N}},y)\,.$

### v4

@@ 4.2 General McKean–Vlasov Setting
In the general setting, under Assumptions 56, and 7, the work of [suzuki2023convergence] provided the first discretization bounds. They impose further assumptions and their resulting complexity bound is rather complicated, but it reads roughly $MN=\widetilde{\mathcal{O}}(\nicefrac{{\operatorname{poly}(\kappa)\,d^{2}}}{{\varepsilon^{4}}})$ for the discretization of ($\mathsf{gMV}_{N}$). Subsequently, [fu2023mean] obtained an improved complexity of $MN=\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{4}d^{3/2}}}{{\varepsilon^{3}}})$ via ULMC in the averaged $\mathsf{TV}$ distance. In comparison, we can improve this complexity guarantee to $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{5/2}d^{3/2}}}{{\varepsilon^{3}}})$, and the guarantee even holds in $\sqrt{\operatorname{\mathsf{KL}}}$ if we combine ULMC with the proximal sampler. It appears that we gain one factor of $\sqrt{\kappa}$ through sharper discretization analysis (via [zhang2023improved], or via the error analysis of the proximal sampler in [altschuler2023faster]), and one factor of $\kappa$ via a sharper propagation of chaos result (Theorem 5).
We also note that the result of [fu2023mean] is based on a kinetic version of the propagation of chaos argument from [Chen+24PoCKinetic], whereas our approach uses the original “non-kinetic” argument from [chen2022uniform] in the form of Theorem 5.
@@
Let us consider the problem of learning a two-layer neural network in the mean-field regime. Let $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ be a function parameterized by $\theta\in\mathbb{R}^{p}$, and for any probability measure $\mu$ over $\mathbb{R}^{p}$, let $f_{\mu}\coloneqq\int f_{\theta}\,\mu(\mathrm{d}\theta)$. For example, in a standard two-layer neural network, we take $\theta=(a,w)\in\mathbb{R}\times\mathbb{R}^{d}$ and $f_{a,w}(x)=a\,\mathsf{ReLU}(\langle w,x\rangle)$. When $\mu=\frac{1}{m}\sum_{j=1}^{m}\delta_{(a_{j},w_{j})}$ is an empirical measure, then $f_{\mu}$ is the function computed by a two-layer neural network with $m$ hidden neurons. In this formulation, however, we can take $\mu$ to be any probability measure, corresponding to the mean-field limit $m\to\infty$ [chizat2018global, mei2018mean, Chi22MFL, RotVan22NNInteracting, sirignano2020mean].
@@
Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ in $\mathbb{R}^{d}\times\mathbb{R}$ and a loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, we can formulate neural network training as the problem of minimizing the loss $\mu\mapsto\sum_{i=1}^{n}\ell(f_{\mu}(x_{i}),y_{i})$. To place this within the general McKean–Vlasov framework, we add two regularization terms: (1) $\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$ corresponds to weight decay; and (2) $\frac{\sigma^{2}}{2}\int\log\mu\,\mathrm{d}\mu$ is entropic regularization. We are now in the setting of §2.1.2, with $\mathcal{F}_{0}(\mu)=\sum_{i=1}^{n}\ell(f_{\mu}(x_{i}),y_{i})$.
To minimize this energy, it is natural to consider the Euler–Maruyama discretization of ($\mathsf{gMV}_{N}$), which corresponds to learning the neural network via noisy GD, and was considered in [suzuki2023convergence]. Recent works [fu2023mean, Chen+24PoCKinetic] also considered the underdamped version of ($\mathsf{gMV}$) and its discretization. Under the assumptions common to those works as well as our own, our results yield improved algorithmic guarantees for this task (see Example 13).
@@
Unfortunately, the assumptions used for the analysis of the general McKean–Vlasov are restrictive and limit the applicability to neural network training. For example, it suffices for $\ell$ to be convex in its first argument (to satisfy Assumption 5), to have two bounded derivatives (w.r.t. its first argument), and for $\theta\mapsto f_{\theta}(x_{i})$ to have two bounded derivatives for each $x_{i}$. The last condition is satisfied, e.g., for $f_{\theta}(x)=\tanh(\langle\theta,x\rangle)$. For a genuinely two-layer example, we can take $f_{\theta}(x)=\tanh(a)\tanh(\langle w,x\rangle)$ for $\theta=(a,w)\in\mathbb{R}\times\mathbb{R}^{d}$. Under these conditions, Assumptions 6 and 8 hold, which in turn furnish log-Sobolev inequalities via Lemmas 6 and 9.
Limitations.
However, we note that there is a substantial limitation of our framework when applied to the mean-field Langevin dynamics. Although we are able to establish a uniform-in-$N$ LSI for the stationary distribution $\mu^{1:N}$ under appropriate assumptions (see Corollary 25 for a precise statement), the dependence of the LSI constant scales poorly (in fact, doubly exponentially) in the problem parameters. To fully benefit from the modularity of our approach, it is desirable to obtain a uniform-in-$N$ LSI with better scaling, and we leave this question open for future research.
@@ 5 Conclusion
We conclude by listing some future directions of study. As discussed in §4.2, our uniform-in-$N$ LSI for the mean-field Langevin dynamics currently scales poorly in the problem parameters, and it is important to improve it. We also believe there is further room for improvement in the propagation of chaos results. For example, can the sharp rate in Theorem 3 be extended to stronger metrics such as Rényi divergences, as well as to situations when the weak interaction condition (Assumption 3) fails, e.g., in the strongly displacement convex case or in the setting of §3.1.2? For the sampling guarantees, the prior works [bou2023nonlinear, suzuki2023convergence] considered different settings, such as potentials satisfying convexity at infinity or the use of stochastic gradients; these extensions are compatible with our approach and could possibly lead to improvements in these cases, as well as others. Finally, consider the case where $\int\nabla W(X_{t}-\cdot)\,\mathrm{d}\pi_{t}$ in ($\mathsf{pMV}$) is replaced with a generic function $\int b_{t}(X_{t},\cdot)\,\mathrm{d}\pi_{t}$, $b_{t}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$. It would be interesting to extend our analysis to this setting, as it arises in many applications [arnaudon2020second].
@@ Acknowledgements
We thank Zhenjie Ren for pointing out an error in a previous draft of this paper, and Daniel Lacker, Atsushi Nitanda, and Taiji Suzuki for important discussions and references. YK was supported in part by NSF awards CCF-2007443 and CCF-2134105. MSZ was supported by NSERC through the CGS-D program. SC was supported by the Eric and Wendy Schmidt Fund at the Institute for Advanced Study. MAE was supported by NSERC Grant [2019-06167] and CIFAR AI Chairs program at the Vector Institute. MBL was supported by NSF grant DMS-2133806.
[…]
@@ Appendix A Control of the Finite-Particle Error
Lemma 14 (Talagrand’s Transport Inequality, [otto2000generalization]).
@@ A.1 LSI Case
$-\frac{\sigma^{2}}{2}\,\nabla_{x^{i}}\log\mu^{1:k}(x^{1:k})$
$\qquad=\nabla V(x^{i})+\frac{1}{N-1}\sum_{\begin{subarray}{c}j=1\\ i\neq j\end{subarray}}^{k}\nabla W(x^{i}-x^{j})+\frac{N-k}{N-1}\operatorname{\mathbb{E}}_{\mu^{k+1|1:k}(\cdot\mid x^{1:k})}\nabla W(x^{i}-\cdot)\,.$
@@ A.1.1 Bounding the Error Terms
$\qquad+(k-1)\,(k-2)\operatorname{\mathbb{E}}\bigl{\langle}\nabla W(X^{1}-X^{2})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot),$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\nabla W(X^{1}-X^{3})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot)\bigr{\rangle}$
@@
$\quad+(k-1)^{2}\operatorname{\mathbb{E}}\bigl{\langle}\nabla W(X^{1}-X^{2})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot),\,\nabla W(X^{1}-X^{3})-\operatorname{\mathbb{E}}_{\pi}\nabla W(X^{1}-\cdot)\bigr{\rangle}\,,$
@@
where the second inequality follows from (TI), and the last one follows from the data-processing inequality for the $\operatorname{\mathsf{KL}}$ divergence. For the second term, the Cauchy–Schwarz inequality leads to
@@ A.2 Strongly Convex Case
The following propagation of chaos argument for the strongly log-concave case is based on [sznitman1991topics]. Let $(X^{1:N}_{t})_{t\geq 0}$ denote the stochastic process following the finite-particle stochastic differential equation ($\mathsf{pMV}_{N}$). Let the corresponding semigroup be denoted ${(\mathcal{T}_{t})}_{t\geq 0}$, defined as follows. For any test function $f:\mathbb{R}^{d\times N}\to\mathbb{R}$,
@@
Proof.  We resort to a coupling argument, noting that $\pi$ is stationary under ($\mathsf{pMV}$). Starting with $\pi^{\otimes N}$, we evolve ${(X^{1:N}_{t})}_{t\geq 0}$ and ${(Y^{1:N}_{t})}_{t\geq 0}$ according to ($\mathsf{pMV}_{N}$) and ($\mathsf{pMV}$) respectively, i.e. $X^{1:N}_{t}\sim\pi^{\otimes N}\mathcal{T}_{t}$ and $Y^{1:N}_{t}\sim\pi^{\otimes N}$. This argument is adapted from the original propagation of chaos proof by [sznitman1991topics].
@@ A.3 General Functional Case
For any measure $\mu$, define its entropy as $\operatorname{\mathsf{ent}}(\mu)=\int\log\mu\,\mathrm{d}\mu$. We now provide a self-contained propagation of chaos argument in the general McKean–Vlasov setting, following [chen2022uniform]. We begin with the following entropy toast inequality, i.e. half of the entropy sandwich inequality from [chen2022uniform].
@@
$\mathcal{E}^{N}(\nu^{1:N})=N\int\mathcal{F}(\rho_{x^{1:N}})\,\nu^{1:N}(\mathrm{d}x^{1:N})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\nu^{1:N})\,.$
@@
$\qquad=-\frac{\sigma^{2}}{2}\operatorname{\mathbb{E}}_{x^{1:N}\sim\nu^{1:N}}\Bigl{[}N\int\log\pi(z)\,(\rho_{x^{1:N}}(\mathrm{d}z)-\pi(\mathrm{d}z))\Bigr{]}+\frac{\sigma^{2}}{2}\,\bigl{(}\operatorname{\mathsf{ent}}(\nu^{1:N})-N\operatorname{\mathsf{ent}}(\pi)\bigr{)}$
@@
$\qquad=-\frac{\sigma^{2}}{2}\int\sum_{i=1}^{N}\log\pi(x^{i})\,\nu^{1:N}(\mathrm{d}x^{1:N})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\nu^{1:N})\,.$
However, this is just $\frac{\sigma^{2}}{2}\operatorname{\mathsf{KL}}(\nu^{1:N}\mathbin{\|}\pi^{\otimes N})$, so we are done. ∎
@@
In order to relate the first term $\mathsf{A}$ to a KL divergence, for each $x^{-i}\in\mathbb{R}^{d\times(N-1)}$ we introduce the probability measure $\tau_{x^{-i}}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ via
@@
$=\log\int\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\delta\mathcal{F}(\pi,z)-\delta\mathcal{F}(\rho_{x^{-i}},z)\bigr{)}\Bigr{)}\,\pi(\mathrm{d}z)+\log Z(\pi)$
@@
$=\frac{\overline{C}_{\mathsf{LSI}}}{\sigma^{2}}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}[\lVert\nabla_{\mathcal{W}_{2}}\mathcal{F}_{0}(\rho_{x^{1:N}},x^{i})-\nabla_{\mathcal{W}_{2}}\mathcal{F}_{0}(\rho_{x^{-i}},x^{i})\rVert^{2}]$
@@
$\zeta_{x^{1:N}}^{i}(y)\coloneqq\delta\mathcal{F}(\rho_{x^{-i}},y)-\delta\mathcal{F}(\rho_{x^{1:N}},y)=\delta\mathcal{F}_{0}(\rho_{x^{-i}},y)-\delta\mathcal{F}_{0}(\rho_{x^{1:N}},y)\,.$